OpenStudy (akohl):

@abhisar Two rocks, each of mass 72 kg, are positioned 95 m away from each other in deep space. What is the magnitude of the gravitational attraction between them? Note: The constant of universal gravity (G) equals 6.674 × 10-11 N · m2/kg2. 3.0 × 10-10 N 3.8 × 10-12 N 3.8 × 10-11 N 9.8 × 10-11 N

OpenStudy (abhisar):

\(\large\sf Force=\frac{G\times M_1\times M_2}{R^2}\)

OpenStudy (akohl):

Thanks I got 3.8*10^-11

OpenStudy (abhisar):

\(\huge\checkmark\)

OpenStudy (akohl):

Posting next question

OpenStudy (abhisar):

YOu should visit this website, it would help u learn better about Openstudy. http://openstudyab.comoj.com

OpenStudy (akohl):

Thanks

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)